OpenStudy (anonymous):

Celeste made a model of her grandfather’s fishing boat using a scale where 2 inches represents 3 feet. Her grandfather’s actual boat is 28 feet long. What is the length of Celeste’s model boat?

OpenStudy (anonymous):

2 inches is 1/6 feet, so let the length of boat is x, then we have this equation: 1/6:3=x:28, cross multiply, we have 3x=28/6, so x =28/18=14/9 feet